Caro Manufacturing has two production departments, Machining and Assembly, and two service departments, Maintenance and Cafeteria. Direct costs for each department and the proportion of service costs used by the various departments for the month of August follow:

Proportion of Services Used by

Department

Direct Costs

Maintenance

Cafeteria

Machining

Assembly

Machining

$

150,000

Assembly

77,000

Maintenance

52,000

0.3

0.4

0.3

Cafeteria

35,000

0.8

0.1

0.1

Required:

Use the step method to allocate the service costs, using the following:

a. The order of allocation starts with Maintenance.

b. The allocations are made in the reverse order (starting with Cafeteria).

  • Required A
  • Required B

The order of allocation starts with Maintenance. (Negative amounts should be indicated by a minus sign. Do not round intermediate calculations.)

To

From

Maintenance

Cafeteria

Machining

Assembly

Service department costs

Maintenance

Cafeteria

Total Costs

$0

$0

$0

$0
Answer #1

Formula : total serive cost × proportion of service used/total service used.

Answer a.

Step method Distribution starts from maintenance service.

Amount of service cost to be distributed.

Cafeteria : $52,000×0.3 /1 = $15,600

Matching : $52,000×0.4/1 = $20,800

Assembly : $52,000×0.3/1 =$15,600.

Now distribution of cafeteria amount to production department in ratio of 1.1

AMOUNT FOR DISTRIBUTION $35,000 + $15,600. = $ 50,600.

Matching : 50,600 ×1/2= $25,300

Assembly : 50,600×1/2=$25,300.

Answer b.

Step method distribution starts from cafeteria

Amount of service cost to be distributed

Maintenance: $ 35,000 × 0.8/1 =$28,000.

Matching: $ 35,000 × 0.1/1 = $3500

Cafeteria: $35,000 ×0.1/1 =$3500.

Now distribution of maintenance to the production department in ratio of 4.3.

AMOUNT FOR DISTRIBUTION $52,000 + $28,000 = $80,000.

Matching: $80,000 × 4/7 = $ 45,714.29

Assembly: $80,000 × 3/7 =$ 34,285.71

For Total distribution

Direct cost + service cost from maintenance and cafeteria respectively.